OpenStudy (anonymous):

Essay, show all work. Consider that the age, x, of a unicorn in human equivalent years can be given by the formula f(x)= -0001518x^4 + 0.067326x^3 - 1.4367x^2 +12.46x +2.914. When a unicorn is 2.5 years old, what is its age in human equivalent years? What about when it is 12 years old? Can someone walk me through this?

OpenStudy (anonymous):

just plug in the age of the unicorn as x. so f(x) = -0001518(2.5)^4 + 0.067326(2.5)^3 - 1.4367(2.5)^2 +12.46(2.5) +2.914 and f(x) = -0001518(12)^4 + 0.067326(12)^3 - 1.4367(12)^2 +12.46(12) +2.914 Then solve the right side of the equation and you have your answer (I recommend a calculator)

OpenStudy (anonymous):

That is suppose to be -0.001518x^4

OpenStudy (anonymous):

ok, the numbers are the x. ok I got it, thank you. I will reply with the answer in a minute. Thank you.

OpenStudy (anonymous):

When a unicorn is 2.5 years old, the age in human equivalent years is: f(x)= -0.001518(2.5)^4 +0.067326(2.5)^3 - 1.4367(2.5)^2 +12.46(2.5) +2.914 f(x)= -0.003795^4 + 0.168315^3 + 3.59175^2 +31.15 +2.914 f(x)= -0.003795^4 + 0.168315^3 + 3.59175^2 +34.064